Metallica Bearings, Inc., is a young start-up company. No dividends will be paid on the stock over the next nine years because the firm needs to plow back its earnings to fuel growth. The company will pay a $3.4 per share dividend 10 years from today and will increase the dividend by 3.24 percent per year thereafter. If the required return on this stock is 7.99 percent, what is the current share price? Omit the $ dollar sign and commas. Do not round intermediate calculations and round your final answer to 2 decimal places. (e.g., 32.16).
Answer #1

Required return on stock (Ke) = 7.99%

Dividend growth rate from year 10 (g) = 3.24%

Dividend in year 10 (D10) = $3.40

Using Gordon's dividend growth model,

Value of share at the end of year 9 = D10 / (ke-g) = 3.40 / ( 7.99%-3.24%) = 71.57895

Therefore current value (price) of share = FV / (1+ke)n = 71.57895 / (1+7.99%)9 = $ 35.84